# Comprehensive survey of hybrid equations of state in neutron star mergers and constraints on the hadron-quark phase transition

### v1

@@ Abstract
We perform an extensive study of equation of state (EoS) models featuring a phase transition from hadronic to deconfined quark matter in neutron star merger simulations. We employ three different hadronic EoSs, a constant speed of sound parametrization for the quark phase and a Maxwell construction to generate a large sample of hybrid EoS models. We systematically vary the onset density and density jump of the phase transition as well as the quark matter stiffness and simulate binary neutron star mergers to infer how the properties of the phase transition affect the gravitational-wave signal. In total we simulate mergers with 245 different hybrid EoS models. In particular, we explore in which scenarios a phase transition would be detectable by a characteristically increased postmerger gravitational-wave frequency compared to an estimate from the inspiral signal assuming a purely hadronic EoS. We find that the density jump at the transition (latent heat) has the largest impact on the gravitational-wave frequencies, while the influence of the stiffness of quark matter is smaller. We quantify which range of phase transition properties would be compatible with a certain magnitude or absence of the gravitational-wave postmerger frequency shift. By means of these dependencies, a future detection will thus directly yield constraints on the allowed features of the hadron-quark phase transition.
@@ I Introduction
At high densities or temperatures matter is expected to undergo a transition from hadronic to deconfined quark matter. Around zero baryon chemical potential, this transition is relatively well explored from collider experiments and lattice quantum chromodynamics (QCD) calculations [1, 2, 3, 4, 5, 6, 7]. At finite baryon chemical potential, the QCD phase diagram is much less understood as the sign problem currently prevents theoretical calculations in this regime [8]. It has long been hypothesized that the transition from hadronic to deconfined quark matter could occur at densities reached in the interiors of neutron stars (NSs) and binary neutron star (BNS) merger remnants [9, 10, 11, 12, 13, 14, 15], see e.g. [16, 17, 18, 19, 20, 21, 22, 23] for reviews. Studying these compact objects and confirming or ruling out the presence of deconfined quark matter therefore provides valuable insights into the QCD phase diagram.
The transition to deconfined quark matter has been studied in the context of BNS merger simulations and core collapse supernovae employing so-called hybrid equation of state (EoS) models, i.e. models with a hadronic and a deconfined quark phase connected by a phase transition [24, 25, 26, 27, 28, 29, 30, 21, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 42, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. These studies are limited to a small set of particular EoS models. Typically, the appearance of deconfined quark matter leads to more compact stars or merger remnants and correspondingly to higher postmerger gravitational-wave (GW) frequencies compared to purely hadronic models. This has been pointed out as a smoking-gun signal to identify the occurrence of the hadron-quark phase transition [29, 31, 34, 37]. Additionally, remnants with hybrid EoS tend to collapse to black holes earlier [30, 36, 38, 54].
@@ II.1 Hybrid equations of state
To construct different hybrid EoSs, we employ a constant speed of sound parametrization of the quark phase following Ref. [57, 56, 58]. This approach is able to reproduce existing microphyiscal models such as NJL based calculations as in [63] or the DD2F-SF models [64] (see Appendix A) very well. We present the resulting equations for pressure, energy and number density in Appendix A.
At zero temperature, we match this quark matter model to one of three hadronic EoSs, namely DD2F [65, 66], DD2 [65, 67] and SFHo [67, 68], using a Maxwell construction [15, 69, 59, 60, 61] (see Sect. II.2 and Ref. [62] for the extension to finite temperatures). This scheme leads to a density jump at the phase transition. We refer to the onset baryon number density of the transition as $n_{\mathrm{on}}$ and the size of the density jump as $\Delta n$. The onset baryon number density of the pure, deconfined quark phase $n_{\mathrm{fin}}$ is then given by $n_{\mathrm{fin}}=n_{\mathrm{on}}+\Delta n$.
@@
Figure 1: Minimum speed of sound squared in the quark phase required to obtain a maximum neutron star mass $M_{\mathrm{max}}\geq 2~{}M_{\odot}$ color-coded as a function of onset density and density jump of the phase transition for hybrid models based on the hadronic SFHo (left panel), DD2F (middle panel) and DD2 (right panel) models. The stiffness of the hadronic models increases from left to right. In each panel the dashed line encloses the region where the requirement $M_{\mathrm{max}}\geq 2~{}M_{\odot}$ cannot be fulfilled. Black symbols mark configurations we simulate in this work with plus signs indicating models that promptly collapse to a black hole at merger for all $c_{s}^{2}\leq 1$. For EoSs with the highest simulated $n_{\mathrm{on}}$ and beyond, the phase transition does not have an impact on the GW signal anymore.
The black symbols mark configurations of $n_{\mathrm{on}}$ and $\Delta n$ we consider in this work. For each pair, we pick different values of $c_{s}^{2}$ ranging from $1$ down to the lowest values indicated in Fig. 1 for the given $\Delta n$ and $n_{\mathrm{on}}$ and simulate mergers of two $1.35~{}M_{\odot}$ stars. Crosses mark configurations allowing for (temporary) stable merger remnants and the occurrence of a strong postmerger GW signal while plus signs depict models where we find prompt black-hole formation at merger even for maximum quark matter stiffness. We also include some configurations with maximum NS masses slightly below $2~{}M_{\odot}$ to cover a broader region of the hybrid model space. We do not consider parametrizations with larger $n_{\mathrm{on}}$ as we find that for those the phase transition no longer impacts the gravitational wave signal since the produced amount of quark matter becomes too small to noticeably affect the structure of the remnant.
@@
Figure 2: Mass-radius curves of cold neutron stars for a selection of hybrid EoSs we employ in this study based on a constant speed of sound model (colored). Each column shows results from hybrid models based on a different underlying hadronic model (black). The top row displays curves of hybrid models with a single onset density of the phase transition, while the bottom row shows hybrid models at a constant speed of sound of $c_{s}^{2}=0.7$ in the quark phase. The stiffness of the hadronic phase increases from left to right.
@@ II.2 Simulation code
We perform merger simulations using a general relativistic smoothed particle hydrodynamics (SPH) code of Refs. [81, 82, 83, 62]. This code solves the Einstein field equations employing the conformal flatness condition [84, 85]. Since the hybrid models we employ (see Sect. II.1) are purely barotropic EoSs [^1] The hadronic models SFHo, DD2F and DD2 are available as fully temperature and composition dependent tables. However, in order to implement thermal effects of quark matter and the temperature dependence of the phase boundaries, we only employ the zero temperature beta-equilibrium slice of these hadronic tables and approximate thermal effects in the hadronic phase. This allows to use the scheme developed in Ref. [62] for modeling thermal effects in hybrid EoSs, which is necessary because the quark matter EoSs of our large sample are only available as barotropes., we additionally include thermal effects using the effPT scheme we presented in Ref. [62]. This scheme extends the commonly used ideal-gas approach [86, 87] by also taking into account the temperature-dependence of the hadron-quark phase transition. It assumes two separated phases of matter connected with a flat, i.e. $\frac{\mathrm{d}P}{\mathrm{d}n}=0$, coexistence region (Maxwell construction) at all temperatures. We briefly summarize the main features of this scheme as well as some slight modifications compared to Ref. [62] in Appendix B.
@@
All simulations are set up using cold, irrotational $1.35-1.35~{}M_{\odot}$ binaries at an initial center-to-center separation of around 38 km. The stars then merge within a few orbits. The effects of neutrinos and magnetic fields are not included. In total, we perform 245 simulations using hybrid models, 71 based the DD2F, 106 based on the DD2 and 68 based on the SFHo EoS.
@@ III Results
with $a=8.463\times 10^{-7}$ kHz, $b=\textendash 2.509\times 10^{-3}$ kHz and $c=4.182$ kHz from Ref. [34] using the tidal deformability of a 1.35 $M_{\odot}$ star from the respective EoS of the inspiraling stars  [^2] Note that we will also refer to this value as $f_{\mathrm{peak,had}}$ if the $n_{\mathrm{on}}$ is below the central density of a $1.35~{}M_{\odot}$ star, i.e. when the inspiraling stars are hybrid stars..
@@
Figure 3: Shift of the dominant postmerger gravitational-wave frequency $f_{\mathrm{peak}}$ in merger simulations using SFHo-based hybrid EoSs compared to $f_{\mathrm{peak}}$ predicted by Eq. (1) based on the tidal deformability of the inspiraling stars (i.e. assuming a hadronic EoS) as a function of the onset density, density jump and stiffness of the quark phase. The dashed line marks the central density of the inspiraling purely hadronic stars, i.e. for SFHo.
In Figs. 34 and 5 we plot $\Delta f_{\mathrm{peak}}$ as a function of $n_{\mathrm{on}}$ and $\Delta n$ for all our simulations that do not result in a prompt collapse at merger (see Appendix C for similar plots using the baryon chemical potential). The coloring indicates $c_{s}^{2}$ in the quark phase. The dashed lines highlight the central density of the inspiraling stars. This means that in all systems with smaller $n_{\mathrm{on}}$ deconfined quark matter is already present prior to the merger while in all systems with larger $n_{\mathrm{on}}$ the two NSs are purely hadronic. We find that $\Delta f_{\mathrm{peak}}$ can vary significantly up to almost 750 Hz depending on the parameters of the hybrid EoS. Generally, $\Delta f_{\mathrm{peak}}$ grows with $\Delta n$ and decreases more weakly with $c_{s}^{2}$. This is understandable, since a larger $\Delta n$ leads to a more compact remnant while a stiffer quark phase reduces the compactness of the quark core compared to a softer quark phase. For very large $\Delta n$ a prompt collapse occurs (no data point shown).
The dependence of $\Delta f_{\mathrm{peak}}$ on the onset density of the phase transition is more complex. At low $n_{\mathrm{on}}$, we find $\Delta f_{\mathrm{peak}}\approx 0$ followed by an increase in $\Delta f_{\mathrm{peak}}$ with $n_{\mathrm{on}}$ eventually reaching a maximum. A further increase in $n_{\mathrm{on}}$ results in decreasing frequency shifts. This behavior is caused by counteracting effects. On the one hand, at low onset densities the inspiraling stars are hybrid stars and the tidal deformability is also affected by the phase transition increasing $f_{\mathrm{peak,had}}$. Hence, the transition is not visible by comparing $\Lambda$ and $f_{\mathrm{peak}}$. Additionally, the quark phase is ypically stiffer than the hadronic phase, as can be seen in Fig. 2 by the larger maximum masses of hybrid EoSs with small density jumps compared to the purely hadronic EoSs. A low onset density, i.e. a large quark core therefore results in relatively low remnant densities despite the softening from the phase transition.
@@
We remark that the magnitude of the frequency shift $\Delta f_{\mathrm{peak}}$ of the hybrid star mergers in Ref. [37] somewhat differ from our values at similar $n_{\mathrm{on}}=1.8\times n_{\mathrm{sat}}$ and $\Delta n=1.45\times n_{\mathrm{sat}}$ with the DD2 base hadronic model. These discrepancies can be attributed to the different finite-temperature behavior of the phase boundary. The transition region in the models employed by Ref. [37] is shifted stronger towards lower densities with increasing temperature than in this study (compare Fig. 10 in the appendix and Fig. 9 in Ref. [64]).
@@
Figure 6: Shift of the dominant postmerger gravitational wave frequency $\Delta f_{\mathrm{peak}}$ in merger simulations using DD2F-based hybrid EoSs compared to $f_{\mathrm{peak}}$ predicted by Eq. (1) based on the tidal deformability of the inspiraling stars as a function of different hybrid EoS properties. (a): $\Delta f_{\mathrm{peak}}$ as a function of $c^{2}$ for $n_{\mathrm{on}}=2.5\times n_{\mathrm{sat}}$. Different colors refer to different $\Delta n$ at the phase transition. (b): $\Delta f_{\mathrm{peak}}$ as a function of $\Delta n$ for $n_{\mathrm{on}}=2.5\times n_{\mathrm{sat}}$. (c): $\Delta f_{\mathrm{peak}}$ as a function of $n_{\mathrm{on}}$ for $\Delta n=0.6\times n_{\mathrm{sat}}$. In the latter two panels different colors refer to results from models with different $c^{2}_{s}$.
@@
For the dependence of $\Delta f$ on $c_{s}^{2}$, a small curvature is visible in Fig. 6. However, we still find reasonable agreement with a linear model. Since $\Delta f_{\mathrm{peak}}$ shows a local maximum with respect to $n_{\mathrm{on}}$, we employ a quadratic dependence in our empirical formula. We find that this approach describes the increase of $\Delta f$ with rising $n_{\mathrm{on}}$ with good accuracy, the generally faster fall-off after the maximum is however not captured so well.
We refrain from exploring more complicated fit formulae that many possibly provide better fits because our data set is still somewhat limited with only three hadronic base models, a simple quark matter parametrization and a single prescribed temperature dependence of the phase boundaries.
@@
We fit this formula with $a,~{}b,~{}c,~{}d$ and $e$ being fit parameters to our dataset for each hadronic model separately using a least squares fit. This is motivated by the expectation that by the time $f_{\mathrm{peak}}$ will be measured, the EoS at lower density and NS mass will be relatively well known e.g. from GW inspiral detections. We restrict the fits to the onset density ranges $n_{\mathrm{on}}\in[2.0\times n_{\mathrm{sat}},~{}3.1\times n_{\mathrm{sat}}]$ for DD2F-based models, $n_{\mathrm{on}}\in[1.6\times n_{\mathrm{sat}},~{}2.1\times n_{\mathrm{sat}}]$ for DD2-based models and $n_{\mathrm{on}}\in[2.0\times n_{\mathrm{sat}},~{}3.5\times n_{\mathrm{sat}}]$ for SFHo-based models, as in this region the amount of quark matter in the remnant is sufficient to influence $f_{\mathrm{peak}}$ and there is a unique dominant postmerger GW frequency visible in the spectrum. At larger $n_{\mathrm{on}}$, $\Delta f_{\mathrm{peak}}$ drops to zero because no or only small amounts of quark matter are present in the merger remnant. We provide an additional argument for these ranges in Appendix. D.
@@
In general, this relation captures the overall behavior of the data relatively well with a mean scatter up to around 60 Hz. However, we also observe deviations from the fit up to 200 Hz, for DD2-based models. The relation we found is hence not very tight for all data points although on average it permits a decent approximation. We omit to provide general formula independent of the hadronic EoS since the range of $n_{\mathrm{on}}$ where $\Delta f_{\mathrm{peak}}$ is substantially larger than zero depends strongly on the properties of hadronic matter.
The relation in Eq. (2) can be used in the future to constrain the properties of a phase transition based on an observed $f_{\mathrm{peak}}$, once sufficiently precise GW measurements are available [93, 94]. First, $f_{\mathrm{peak}}$ together with the inferred tidal deformability can be compared to (1) (or an equivalent formula if the binary mass differs significantly from $2\times 1.35~{}\mathrm{M}_{\odot}$) to obtain $f_{\mathrm{peak,had}}$ and hence $\Delta f_{\mathrm{peak}}$. If $\Delta f_{\mathrm{peak}}$ is at least $\sim$200 Hz, this indicates the presence of a phase transition in the merger remnant. Then, Eq. (2) can be inverted to obtain constraints on $c_{s}$, $n_{\mathrm{on}}$ or $\Delta n$.
@@
Hypothetically assuming that a future detection would yield a finite $\Delta f_{\mathrm{peak}}$, this would constrain the allowed values of $n_{\mathrm{on}}$ and $\Delta n$ to a relatively narrow range depending on the hadronic EoS below $n_{\mathrm{on}}$. We remark that once sufficiently precise GW measurements to infer $f_{\mathrm{peak}}$ are available, the cold EoS at lower densities may be relatively well known from GW inspiral signals, but we also refer to the limitations of our current study briefly addressed in the next section, which may imply a moderate broadening of the parameter ranges discussed above, e.g., by other hadronic base models, different phase boundaries at finite temperature or other quark matter models.

### v2

@@ Abstract
We perform an extensive study of equation of state (EoS) models featuring a phase transition from hadronic to quark matter in neutron star merger simulations. We employ three different hadronic EoSs, a constant speed of sound parameterization for the quark phase and a Maxwell construction to generate a large sample of hybrid EoS models. We systematically vary the onset density and density jump of the phase transition as well as the quark matter stiffness and simulate binary neutron star mergers to infer how the properties of the phase transition affect the gravitational-wave signal. In total we simulate mergers with 245 different hybrid EoS models. In particular, we explore in which scenarios a phase transition would be detectable by a characteristically increased postmerger gravitational-wave frequency compared to an estimate from the inspiral signal assuming a purely hadronic EoS. We find that the density jump at the transition (latent heat) has the largest impact on the gravitational-wave frequencies, while the influence of the stiffness of quark matter is smaller. We quantify which range of phase transition properties would be compatible with a certain magnitude or absence of the gravitational-wave postmerger frequency shift. By means of these dependencies, a future detection will thus directly yield constraints on the allowed features of the hadron-quark phase transition.
@@ I Introduction
At high densities or temperatures matter is expected to undergo a transition from hadronic to quark matter. Around zero baryon chemical potential, this transition is relatively well explored from collider experiments and lattice quantum chromodynamics (QCD) calculations [1, 2, 3, 4, 5, 6, 7]. At finite baryon chemical potential, the QCD phase diagram is much less understood as the sign problem currently prevents theoretical calculations in this regime [8]. It has long been hypothesized that the transition from hadronic to deconfined quark matter could occur at densities reached in the interiors of neutron stars (NSs) and binary neutron star (BNS) merger remnants [9, 10, 11, 12, 13, 14, 15], see e.g. [16, 17, 18, 19, 20, 21, 22, 23] for reviews. Studying these compact objects and confirming or ruling out the presence of deconfined quark matter therefore provides valuable insights into the QCD phase diagram.
The transition to deconfined quark matter has been studied in the context of BNS merger simulations and core collapse supernovae employing so-called hybrid equation of state (EoS) models, i.e. models with a hadronic and a deconfined quark phase connected by a phase transition [24, 25, 26, 27, 28, 29, 30, 21, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 42, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. These studies are limited to a small set of particular EoS models. Typically, the appearance of deconfined quark matter leads to more compact stars or merger remnants and correspondingly to higher postmerger gravitational-wave (GW) frequencies compared to purely hadronic models. This has been pointed out as a smoking-gun signal to identify the occurrence of the hadron-quark phase transition [29, 31, 34, 37]. Additionally, remnants with hybrid EoS tend to collapse to black holes earlier [30, 36, 38, 54].
@@ II.1 Hybrid equations of state
To construct different hybrid EoSs, we employ a constant speed of sound parametrization of the quark phase following Ref. [59, 58, 60]. This approach is able to reproduce existing microphyiscal models such as NJL based calculations as in [65] or the DD2F-SF models [66] (see Appendix A) very well. We present the resulting equations for pressure, energy and number density in Appendix A. We emphasize that our approach is independent of the microphysics of the phase transition and of the quark matter EoS, and we do not explicitly distinguish between the chiral phase transition and the onset of deconfinement. For the purpose of this study, the model independence is a particular advantage because it allows to tune the phase transition properties and to build a large set of models, which was not possible by employing full microphysical EoS tables that are not available in large numbers and do not offer the flexibility to adapt the parameters of the phase transition. We note that thermal effects of the EoS are included by an effective scheme with the most prominent impact emerging from the temperature dependence of the phase boundaries, which should be considered as additional degree of freedom of the EoS (see Sect. II.2, App. B and [64]).
At zero temperature, we match the constant-speed-of-sound quark matter model to one of three hadronic EoSs, namely DD2F [67, 68], DD2 [67, 69] and SFHo [69, 70], using a Maxwell construction [15, 71, 61, 62, 63]. This scheme leads to a density jump at the phase transition. We refer to the onset baryon number density of the transition as $n_{\mathrm{on}}$ and the size of the density jump as $\Delta n$. The onset baryon number density of the pure, deconfined quark phase $n_{\mathrm{fin}}$ is then given by $n_{\mathrm{fin}}=n_{\mathrm{on}}+\Delta n$.
@@
(a) SFHo
(b) DD2F
(c) DD2
Figure 1: Minimum speed of sound squared in the quark phase required to obtain a maximum neutron star mass $M_{\mathrm{max}}\geq 2~M_{\odot}$ color-coded as a function of onset density and density jump of the phase transition for hybrid models based on the hadronic SFHo (left panel), DD2F (middle panel) and DD2 (right panel) models. The stiffness of the hadronic models increases from left to right. In each panel the dashed line encloses the region where the requirement $M_{\mathrm{max}}\geq 2~M_{\odot}$ cannot be fulfilled. Black symbols mark EoS models for which we perform simulations of binary NS mergers in this work with plus signs indicating models that promptly collapse to a black hole at merger for all $c_{s}^{2}\leq 1$. For EoSs with the highest simulated $n_{\mathrm{on}}$ and beyond, the phase transition does not have an impact on the GW signal anymore.
The black symbols mark configurations of $n_{\mathrm{on}}$ and $\Delta n$, i.e. EoS models, for which we perform NS merger simulations in this work. For each pair, we pick different values of $c_{s}^{2}$ ranging from $1$ down to the lowest values indicated in Fig. 1 for the given $\Delta n$ and $n_{\mathrm{on}}$ and simulate mergers of two $1.35~M_{\odot}$ stars. Crosses mark configurations allowing for (temporary) stable merger remnants and the occurrence of a strong postmerger GW signal while plus signs depict models where we find prompt black-hole formation at merger even for maximum quark matter stiffness. We also include some configurations with maximum NS masses slightly below $2~M_{\odot}$ to cover a broader region of the hybrid model space. We do not consider parametrizations with larger $n_{\mathrm{on}}$ as we find that for those the phase transition no longer impacts the gravitational wave signal since the produced amount of quark matter becomes too small to noticeably affect the structure of the remnant.
@@
Figure 2: Mass-radius curves of cold neutron stars for a selection of hybrid EoSs we employ in this study based on a constant speed of sound model (colored). Each column shows results from hybrid models based on a different underlying hadronic model (black). The top row displays curves of hybrid models with a single onset density of the phase transition of $2\times n_{\mathrm{sat}}$ (SFHO), $2\times n_{\mathrm{sat}}$ (DD2F) and $1.8\times n_{\mathrm{sat}}$ (DD2). The numbers indicate the density jump $\Delta n$ at the phase transition in units of nuclear saturation density. The bottom row shows hybrid models at a constant speed of sound of $c_{s}^{2}=0.7$ in the quark phase. The stiffness of the hadronic phase increases from left to right.
@@ II.2 Simulation code
We perform merger simulations using the general relativistic smoothed particle hydrodynamics (SPH) code of Refs. [83, 84, 85, 64]. This code solves the Einstein field equations employing the conformal flatness condition [86, 87]. The code employs a cubic spline kernel with an adaptive smoothing length [88]. A time-dependent artificial viscosity scheme is included with an additional switch to reduce viscosity in pure shear flows [89, 90, 91, 92]. Since the hybrid models we employ (see Sect. II.1) are purely barotropic EoSs [^1] The hadronic models SFHo, DD2F and DD2 are available as fully temperature and composition dependent tables. However, in order to implement thermal effects of quark matter and the temperature dependence of the phase boundaries, we only employ the zero temperature beta-equilibrium slice of these hadronic tables and approximate thermal effects in the hadronic phase. This allows to use the scheme developed in Ref. [64] for modeling thermal effects in hybrid EoSs, which is necessary because the quark matter EoSs of our large sample are only available as barotropes., we additionally include thermal effects using the effPT scheme we presented in Ref. [64]. This scheme extends the commonly used ideal-gas approach [93, 94] by also taking into account the temperature-dependence of the hadron-quark phase transition. It assumes two separated phases of matter connected with a flat, i.e. $\frac{\mathrm{d}P}{\mathrm{d}n}=0$, coexistence region (Maxwell construction) at all temperatures. We briefly summarize the main features of this scheme as well as some slight modifications compared to Ref. [64] in Appendix B.
@@
The comparison in [64] shows a very good agreement between merger simulations with fully temperature and composition dependent EoS models and calculations with the effective scheme (see Tab. II in [64] with postmerger GW frequencies coinciding to within a few 10 Hz.).
All simulations are set up using cold, irrotational $1.35-1.35~M_{\odot}$ binaries at an initial center-to-center separation of around 38 km. The stars then merge within a few orbits. Each star is modeled by about 150,000 SPH particles and the number of SPH particle neighbors is kept at about 100 throughout the simulation. For an SPH particle number in this range, the impact of resolution on the dominant GW frequency is small (a few 10 Hz, e.g. [95, 96]). The effects of neutrinos and magnetic fields are not included. In total, we perform 245 simulations using hybrid models, 71 based the DD2F, 106 based on the DD2 and 68 based on the SFHo EoS [^2] The specific models and their simulation results are available in the Supplemental Material [97]..
@@ III Results
with $a=8.463\times 10^{-7}$ kHz, $b=\textendash 2.509\times 10^{-3}$ kHz and $c=4.182$ kHz from Ref. [34] using the tidal deformability of a 1.35 $M_{\odot}$ star from the respective EoS of the inspiraling stars  [^3] Note that we will also refer to this value as $f_{\mathrm{peak,had}}$ if the $n_{\mathrm{on}}$ is below the central density of a $1.35~M_{\odot}$ star, i.e. when the inspiraling stars are hybrid stars. $f_{\mathrm{peak,had}}$ is the fiducial value one expects for a given tidal deformability if there is no quark matter present at all..
@@
Figure 3: Shift of the dominant postmerger gravitational-wave frequency $f_{\mathrm{peak}}$ in merger simulations using SFHo-based hybrid EoSs compared to $f_{\mathrm{peak}}$ predicted by Eq. (1) based on the tidal deformability of the inspiraling stars (i.e. assuming a hadronic EoS) as a function of the onset density, density jump and stiffness of the quark phase. The dashed line marks the central density of the inspiraling purely hadronic stars, i.e. for SFHo. Hence, models at smaller values of $n_{\mathrm{on}}$ contain quark matter before merging, whereas models to the right of the dashed line are purely hadronic during the inspiral.
In Figs. 34 and 5 we plot $\Delta f_{\mathrm{peak}}$ as a function of $n_{\mathrm{on}}$ and $\Delta n$ for all our simulations that do not result in a prompt collapse at merger (see Appendix C for plots using the baryon chemical potential showing similar functional dependencies). The coloring indicates $c_{s}^{2}$ in the quark phase. The dashed lines highlight the central density of the inspiraling stars. This means that in all systems with smaller $n_{\mathrm{on}}$ deconfined quark matter is already present prior to the merger while in all systems with larger $n_{\mathrm{on}}$ the two NSs are purely hadronic. We find that $\Delta f_{\mathrm{peak}}$ can vary significantly up to almost 750 Hz depending on the parameters of the hybrid EoS. Generally, $\Delta f_{\mathrm{peak}}$ grows with $\Delta n$ and decreases more weakly with $c_{s}^{2}$. This is understandable, since a larger $\Delta n$ leads to a more compact remnant while a stiffer quark phase reduces the compactness of the quark core compared to a softer quark phase. For very large $\Delta n$ a prompt collapse occurs (no data point shown).
The dependence of $\Delta f_{\mathrm{peak}}$ on the onset density of the phase transition is more complex. At low $n_{\mathrm{on}}$, we find $\Delta f_{\mathrm{peak}}\approx 0$ followed by an increase in $\Delta f_{\mathrm{peak}}$ with $n_{\mathrm{on}}$ eventually reaching a maximum. A further increase in $n_{\mathrm{on}}$ results in decreasing frequency shifts. This behavior is caused by counteracting effects. On the one hand, at low onset densities the inspiraling stars are hybrid stars and the tidal deformability is also affected by the phase transition increasing $f_{\mathrm{peak,had}}$. Hence, the transition is not visible by comparing $\Lambda$ and $f_{\mathrm{peak}}$. Additionally, the quark phase is typically stiffer than the hadronic phase, as can be seen in Fig. 2 by the larger maximum masses of hybrid EoSs with small density jumps compared to the purely hadronic EoSs. A low onset density, i.e. a large quark core therefore results in relatively low remnant densities despite the softening from the phase transition. For some models with low $n_{\mathrm{on}}$, we find a negative frequency shift, i.e. the hybrid model with quark matter leads to a decrease of the dominant postmerger GW frequency compared to the one from the corresponding purely nucleonic model (cf. [103]).
@@
We remark that the magnitude of the frequency shift $\Delta f_{\mathrm{peak}}$ of the hybrid star mergers in Ref. [37] somewhat differ from our values at similar $n_{\mathrm{on}}=1.8\times n_{\mathrm{sat}}$ and $\Delta n=1.45\times n_{\mathrm{sat}}$ with the DD2 base hadronic model. These discrepancies can be attributed to the different finite-temperature behavior of the phase boundary adopted in this work as compared to [37]. The transition region in the models employed by Ref. [37] is shifted stronger towards lower densities with increasing temperature than in this study (compare Fig. 10 in the appendix and Fig. 9 in Ref. [66]).
@@
Figure 6: Shift of the dominant postmerger gravitational wave frequency $\Delta f_{\mathrm{peak}}$ in merger simulations using DD2F-based hybrid EoSs compared to $f_{\mathrm{peak}}$ predicted by Eq. (1) based on the tidal deformability of the inspiraling stars as a function of different hybrid EoS properties. (a): $\Delta f_{\mathrm{peak}}$ as a function of $c_{s}^{2}$ for $n_{\mathrm{on}}=2.5\times n_{\mathrm{sat}}$. Different colors refer to different $\Delta n$ at the phase transition. (b): $\Delta f_{\mathrm{peak}}$ as a function of $\Delta n$ for $n_{\mathrm{on}}=2.5\times n_{\mathrm{sat}}$. (c): $\Delta f_{\mathrm{peak}}$ as a function of $n_{\mathrm{on}}$ for $\Delta n=0.6\times n_{\mathrm{sat}}$. In the latter two panels different colors refer to results from models with different $c^{2}_{s}$.
@@
For the dependence of $\Delta f$ on $c_{s}^{2}$, a small curvature is visible in Fig. 6. However, we still find reasonable agreement with a linear model (better than 100 Hz). Since $\Delta f_{\mathrm{peak}}$ shows a local maximum with respect to $n_{\mathrm{on}}$, we employ a quadratic dependence in our empirical formula. We find that this approach describes the increase of $\Delta f$ with rising $n_{\mathrm{on}}$ with good accuracy, the generally faster fall-off after the maximum is however not captured so well.
We refrain from exploring more complicated fit formulae that may possibly provide better fits because our data set is still somewhat limited with only three hadronic base models, a simple quark matter parametrization and a single prescribed temperature dependence of the phase boundaries.
@@
We fit this formula with a,b,c,da,~b,~c,~d and $e$ being fit parameters to our dataset for each hadronic model separately using a least squares fit. This is motivated by the expectation that by the time $f_{\mathrm{peak}}$ will be measured, the EoS at lower density and NS mass will be relatively well known e.g. from GW inspiral detections. We omit to provide a general formula independent of the hadronic EoS since the range of $n_{\mathrm{on}}$ where $\Delta f_{\mathrm{peak}}$ is substantially larger than zero depends strongly on the properties of hadronic matter. We restrict the fits to the onset density ranges $n_{\mathrm{on}}\in[2.0\times n_{\mathrm{sat}},~3.1\times n_{\mathrm{sat}}]$ for DD2F-based models, $n_{\mathrm{on}}\in[1.6\times n_{\mathrm{sat}},~2.1\times n_{\mathrm{sat}}]$ for DD2-based models and $n_{\mathrm{on}}\in[2.0\times n_{\mathrm{sat}},~3.5\times n_{\mathrm{sat}}]$ for SFHo-based models, as in this region the amount of quark matter in the remnant is sufficient to influence $f_{\mathrm{peak}}$ and there is a unique dominant postmerger GW frequency visible in the spectrum. At larger $n_{\mathrm{on}}$, $\Delta f_{\mathrm{peak}}$ drops to zero because no or only small amounts of quark matter are present in the merger remnant. We provide an additional argument for these ranges in Appendix. D.
@@
In general, this relation captures the overall behavior of the data relatively well with a mean scatter up to around 60 Hz. However, we also observe deviations from the fit up to 200 Hz, for DD2-based models. The relation we found is hence not very tight for all data points although on average it permits a decent approximation.
The relation in Eq. (2) can be used in the future to constrain the properties of a phase transition based on an observed $f_{\mathrm{peak}}$, once sufficiently precise GW measurements are available [104, 105]. In this regard we mention that the postmerger signal-to-noise ratios (SNRs) of GW signals from simulations with the hybrid EoSs are very comparable to those of purely nucleonic models (for noise curves of a single Advanced LIGO detector at design sensitivity [106] the SNRs vary between 0.35 and 2.17 with an average of 1.36 across all hybrid models compared to 1.5295, 1.497, and 2.0205 for the nucleonic EoSs SFHO, DD2F and DD2 at a distance of 40 Mpc along the polar direction [^4] Note that the SNR increases if a network of detectors is used.). For a signal comparably loud as that of GW170817, an improvement of the sensitivity by factors of a few compared to Advanced LIGO’s design sensitivity is required to extract $f_{\mathrm{peak}}$ [104, 107, 105]. First, $f_{\mathrm{peak}}$ together with the inferred tidal deformability can be compared to (1) (or an equivalent formula if the binary mass differs significantly from $2\times 1.35~\mathrm{M}_{\odot}$; see e.g. [34] for typical variations with $M_{\mathrm{tot}}$ and $q$, which are for instance of the order of a few 10 Hz for changes of $10^{-2}~M_{\odot}$ in $M_{\mathrm{tot}}$) to obtain $f_{\mathrm{peak,had}}$ and hence $\Delta f_{\mathrm{peak}}$. If $\Delta f_{\mathrm{peak}}$ is at least $\sim$200 Hz, this indicates the presence of a phase transition in the merger remnant. Then, Eq. (2) can be inverted to obtain constraints on $c_{s}$, $n_{\mathrm{on}}$ or $\Delta n$. Employing a refined fit in a smaller parameter range, comparing directly to the data, and computing additional models in a preselected range may be advantageous in this respect.
@@
Hypothetically assuming that a future detection would yield a finite $\Delta f_{\mathrm{peak}}$, this would constrain the allowed values of $n_{\mathrm{on}}$ and $\Delta n$ to a relatively narrow range depending on the hadronic EoS below $n_{\mathrm{on}}$ (see Figs. 34 and 5 for the range of phase transition properties which yield a finite $\Delta f_{\mathrm{peak}}$).. We remark that once sufficiently precise GW measurements to infer $f_{\mathrm{peak}}$ are available, the cold EoS at lower densities may be relatively well known from GW inspiral signals, but we also refer to the limitations of our current study briefly addressed in the next section, which may imply a moderate broadening of the parameter ranges discussed above, e.g., by other hadronic base models, different phase boundaries at finite temperature or other quark matter models.